# Causal Inference for a Hidden Treatment

### v1

@@ 3 Identification of ATE
This bias will in general not be null, unless the ATE of $A^{*}$ on $Y$ is actually null, or if misclassification error is absent. Although unlikely, it may also happen that under certain exceptional laws, the conditional bias $\delta(X)$ given $X$ itself is not null, with opposite signs over different parts of the support of $X$, such that however it averages over $X$ to equal zero.
@@
We implicitly assume that $A\neq A^{*}$ and $Z\neq A^{*}$, a necessary implication of misclassification error. If $A$ equals $A^{*}$, the matrix $P(Y,Z,A=1)$, which will be introduced shortly, would not be of full rank. Therefore, a non-zero determinant of $P(Y,Z,A=1)$ indicates that $A\neq A^{*}$. A similar argument applies to $Z$. This re-iterates the fact that we are primarily concerned with the most challenging settings where misclassification error cannot be ruled out empirically with certainty.
@@
which as noted in Hu (2008) is the canonical form of the eigendecomposition of the matrix on the left-hand side. Since the right-hand side is observed, we can solve for the eigenvalues and eigenvectors, which correspond to the diagonal elements of $P_{D}(A=1\mid A^{*})$ and the columns of $P(Y\mid A^{*})$, respectively. While eigenvectors are identified up to scale, the requirement that columns of $P(Y\mid A^{*})$ sum to 1 fixes it. The remaining challenge lies in correctly matching each eigenvalue with its respective diagonal element in $P_{D}(A=1\mid A^{*})$. To resolve this, we propose the following assumption which encodes four alternative ways in which the surrogate $A$ carries sufficient information about $A^{*}$ to resolve the remaining identification gap.
@@ 4. Condition 4
Assumption 4 clarifies the different roles of $A$ and $Z$, although they appear symmetric in Figure 1. While both the proxy $Z$ and the surrogate $A$ need to satisfy Assumptions 2 and 3, an extra Assumption 4 is imposed on the surrogate. This assumption essentially suggests that $A$, as a surrogate, should provide more information about $A^{*}$ than a random guess. The condition can reasonably be expected to hold if $A$ is a useful measurement of $A^{*}$ in the sense that it reflects a genuine albeit imperfect effort to measure the latter. Condition 1 is equivalent to $\textrm{pr}(A=1\mid A^{*}=0,X=x)<\textrm{pr}(A=1\mid A^{*}=1,X=x)$. Similarly, Condition 2 can be written as $\textrm{pr}(A^{*}=1\mid A=0,X=x)<\textrm{pr}(A^{*}=1\mid A=1,X=x)$. It is noteworthy that Condition 1 and Condition 2 are effectively of similar but symmetric nature. Condition 3 indicates the precision $\textrm{pr}(A^{*}=1\mid A=1,X=x)>0.5$, while Condition 4 states that the false omission rate $\textrm{pr}(A^{*}=1\mid A=0,X=x)<0.5$. These conditions are intuitively reasonable in many scenarios. For example, consider $A^{*}$ as the actual COVID-19 vaccination status of a patient, and $A$ as whether the patient was vaccinated at a specific hospital. If hospital records show a positive vaccination status, it is must be that the patient was truly vaccinated, making $\textrm{pr}(A^{*}=1\mid A=1,X=x)=1$, thereby fulfilling Condition 2 (and Condition 1) of Assumption 4. If one can reasonably assume that the misclassification error is balanced across covariates, such that $X\!\perp\!\!\!\perp A\mid A^{*}$, i.e., the surrogate $A$ only depends on $A^{*}$, then one can safely omit $X$ from Assumption 4. We note that Hu (2008) considers alternatives to Conditions 1-4.
The informal proof given above is formally stated in Theorem 1 below in a more general settings where $Y$ can be either binary, a count, or a continuous outcome, and $X$ is possibly vector-valued. For the proof of this theorem, please refer to the Supplementary Materials.
@@ 4 Semiparametric theory
$\Lambda^{F,\bot}=\left\{K=k\left(Y,A,A^{*},Z,X\right):E\left(K|Y,A^{*},X\right)=E\left(K|A,A^{*},X\right)=E\left(K|Z,A^{\ast},X\right)=0\right\}\cap L_{0}^{2,F},$ (10)
@@
The derivation of (10) and (11) is deferred to the Supplementary Materials. In the following, we will make reference to the oracle nonparametric model $\mathcal{M^{*}}_{np}$ as the collection of oracle laws compatible with the DAG given in Figure 2 which places no restriction on the joint distribution of $(A^{*},Y,Z,A,X)$. We then have the following result:
@@
The efficient influence function of $\psi_{a^{*}}$ in the oracle hidden treatment model $\mathcal{M^{*}}$ is equal to the efficient influence function in the oracle nonparametric model $\mathcal{M^{*}}_{np}$.
The proof of Lemma 1 is relegated to the Supplementary Materials.
@@
So far, we have characterized the tangent space, the set of influence functions and the efficient influence function for $\psi_{a^{*}}$ for $\mathcal{M^{*}}$. We now turn our attention to the observed data model $\mathcal{M}$. The following result summarizes our characterization of the tangent space of the observed data hidden treatment model $\Lambda$, and its ortho-complement, obtained by carefully applying a certain coarsening operator which maps the tangent space for the full data model, $\Lambda^{F}$, onto the $L^{2}_{0}(Y,A,Z,X)$ space.
@@
$\Lambda^{\bot}=\{L=l(Y,A,Z,X):E(L\mid Y,A^{*},X)=E(L\mid A,A^{*},X)=E(L\mid Z,A^{*},X)=0\}\cap L^{2}_{0}.$
Suppose $E(Z\mid A^{*}=a^{*},X)\neq E(A\mid A^{*}=a^{*},X),a^{*}=0,1$, almost surely, then $\Lambda^{\bot}$ can be equivalently expressed as
@@
This implies that $h_{1}$ and $h_{2}$ belong to the null space of the $k\times 2$ matrix $M_{Y\mid A^{*},X}$. Under Assumption 2, the null space has a dimension of $k-2$ when $k\geq 2$. When $Y$ is binary, i.e., $k=2$, $M_{Y\mid A^{*},X}$ is full rank, leading to the unique solution of (13): $h_{j}(Y,X)=0,j=1,2$. This is further detailed in Corollary 1. When $Y$ is categorical, see the Supplementary Materials for a closed-form characterization of $\Lambda^{\bot}$.
@@
Suppose that $A^{*},A,Z,Y$ are all binary variables, then the tangent space of $\mathcal{M}$ is $L^{2}_{0}$, i.e., $\Lambda=\{L=l(Y,A,Z,X)\in L^{2}_{0}\}$.
Remark 2.
For continuous $Y$, solving $h(Y,X)$ from $E\{h(Y,X)\mid A^{*},X\}=0$ can be a challenging task which we leave for future work.
Having characterized the observed data tangent space, we note that the set of all observed data influence functions for the functional $\psi_{a^{*}}$ in the hidden treatment model can be written as $\phi+\Lambda^{\bot}$ for any influence function $\phi$. Therefore, it suffice to identify a single influence function $\phi$. To accomplish this, we establish a formal relationship between influence functions in the observed and oracle data models. As we have characterized the influence functions for the oracle model, we may leverage this relationship to derive the corresponding set of influence functions in the observed hidden treatment model. The following result provides such a characterization of observed data influence functions of $\psi_{a^{*}}$.
@@
The efficient influence function $\phi_{a^{*}}^{eff}$ is given by the orthogonal projection of $\phi_{a^{*}}$ onto the observed data tangent space of the hidden treatment model, i.e. $\phi_{a^{*}}^{eff}=\Pi(\phi_{a^{*}}|\Lambda),$ where $\Pi$ is the orthogonal projection operator. In the case of binary $Y$, (3) is the efficient influence function for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$.
The theorem provides interesting insight about the relationship between an oracle influence function and the corresponding observed data influence function. Specifically, comparing (3) with the full data influence function (2), one may note that the former is obtained by substituting the indicator function $\mathbb{I}_{a^{*}}(A^{*})$ in the latter with the following expression:
@@
$=$ $\mathbb{I}_{a^{*}}(A^{*}).$
This provides an elegant link between (2) and (3). In the full data influence function (2), the unobserved variable $A^{*}$ only appears within the indicator function $\mathbb{I}_{a^{*}}(A^{*})$. In the observed data influence function (3), $\mathbb{I}_{a^{*}}(A^{*})$ is replaced by an expression of observed variables (16), whose expectaction conditional on $(A^{*},X)$, is equal to $\mathbb{I}_{a^{*}}(A^{*})$.
@@ 5 Semiparametric estimation and inference
Typically, nuisance functions are estimated in a initial step from the observed data. However, because the nuisance functions in (17) involve the unobserved variable $A^{*}$, estimation of these functions is not straightforward, particularly using nonparametric estimators. A possible approach is to first estimate the observed data distribution, and to then compute nuisance functions following a standard plug-in principle. For a detailed discussion of this approach, please see the Supplementary Materials. A notable challenge of the approach is its strong dependence on the precision with which the observed data distribution $\textrm{pr}(Y,A,Z\mid X)$ can be estimated. This dependence arises because the nuisance functions are actually solutions to polynomial equations that are smooth in the observed data law $\textrm{pr}(Y,A,Z\mid X)$. This dependence renders the estimated nuisance functions excessively sensitive to the uncertainty in $\widehat{\textrm{pr}}(Y,A,Z\mid X)$, which makes it challenging to obtain estimates of the nuisance functions that are guaranteed to remain in the model. To address this, we propose an algorithm that circumvents the need to obtain an initial estimator of the observed data law. The algorithm is a semiparametric EM algorithm, inspired by the nonparametric EM algorithm proposed in Benaglia et al. (2009), see Algorithm 1 for implementation details.
@@
In practice, estimation processes are usually combined with cross-fitting to prevent over-fitting. This is achieved by using separate parts of the dataset for estimating nuisance functions and calculating the empirical mean in (17). Our algorithm is implemented using a modified version of cross-fitting. This approach slightly deviates from the standard cross-fitting procedure, as the training and evaluation sets are iterated simultaneously. Specifically, during each M-step, models $\widehat{E}(p^{(t)}\mid X),\widehat{E}(p^{(t)}\mid Y,X),\widehat{E}(p^{(t)}\mid A,X),\widehat{E}(p^{(t)}\mid Z,X)$ are developed using the training set. These models are then applied to both the training set and the evaluation set to obtain $\widetilde{\textrm{pr}}(A^{*}=1\mid X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Y_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid A_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Z_{i},X_{i})$ for $i=1,\ldots,n$. See Algorithm S1 in the Supplementary Materials for a detailed description. It is important to note that in this process, the models evaluating the nuisances are independent of the evaluation data, making it a valid form of cross-fitting.
@@
Suppose that the nuisance functions are estimated from a separate independent sample, e.g., by cross-fitting. Moreover, suppose that Assumptions 2-5 hold, and
@@
Remarkably, Theorem 4 does not require nuisance functions to converge at root-$n$ rates as it suffices for their pairwise product of convergence rates to be faster than root-$n$, and for the converge rate of $\widehat{E}(A\mid a^{*}_{j},X)$ rates raised to the $\alpha$ power to likewise be faster than root-$n$, where $\alpha$ is a user-specified positive integer. The choice of $\alpha$ appears to offer a cost-free advantage; however, there is indeed a trade-off between $\alpha$ and the margin constant $\epsilon_{1}$ in Assumption 5. If $\epsilon_{1}$ tends to zero, $\alpha$ must also tend to zero, otherwise the constant term $\epsilon_{1}^{-\alpha}$ scaling $\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty}$ will become unbounded (see the Proof of Theorem 4 in the Supplementary Materials.

### v2

@@ 3 Identification of ATE
This bias will generally be non-null, unless the ATE of $A^{*}$ on $Y$ is actually null, or if misclassification error is absent. Although unlikely, it may also happen that under certain exceptional laws, the conditional bias $\delta(X)$ given $X$ itself is not null, with opposite signs over different parts of the support of $X$, such that however it averages over $X$ to equal zero.
@@
We implicitly assume that $A\neq A^{*}$ and $Z\neq A^{*}$, a necessary implication of misclassification error. If $A$ equals $A^{*}$, the matrix $P(Y,Z,A=1)$ which will be introduced shortly, would not be of full rank. Therefore, a non-zero determinant of $P(Y,Z,A=1)$ indicates that $A\neq A^{*}$. A similar argument applies to $Z$. This re-iterates the fact that we are primarily concerned with the most challenging settings where misclassification error cannot be ruled out empirically.
@@
which as noted in Hu (2008) is the canonical form of the eigendecomposition of the matrix on the left-hand side. Since the right-hand side is observed, we can solve for the eigenvalues and eigenvectors, which correspond to the diagonal elements of $P_{D}(A=1\mid A^{*})$ and the columns of $P(Y\mid A^{*})$, respectively. While eigenvectors are identified up to a scale, the requirement that columns of $P(Y\mid A^{*})$ sum to 1 fixes it. The remaining challenge lies in correctly matching each eigenvalue with its respective diagonal element in $P_{D}(A=1\mid A^{*})$. To resolve this, we propose the following assumption which encodes four alternative ways in which the surrogate $A$ carries sufficient information about $A^{*}$ to resolve the remaining identification gap.
@@ 4. Condition 4
Assumption 4 clarifies the different roles of $A$ and $Z$, although they appear symmetric in Figure 1. While both the proxy $Z$ and the surrogate $A$ need to satisfy Assumptions 2 and 3, an extra Assumption 4 is imposed on the surrogate. This assumption essentially suggests that $A$, as a surrogate, should provide more information about $A^{*}$ than a random guess. The condition can reasonably be expected to hold if $A$ is a useful measurement of $A^{*}$ in the sense that it reflects a genuine albeit imperfect effort to measure the latter. Condition 1 is equivalent to $\textrm{pr}(A=1\mid A^{*}=0,X=x)<\textrm{pr}(A=1\mid A^{*}=1,X=x)$. Similarly, Condition 2 can be written as $\textrm{pr}(A^{*}=1\mid A=0,X=x)<\textrm{pr}(A^{*}=1\mid A=1,X=x)$. It is noteworthy that Condition 1 and Condition 2 are effectively of similar but symmetric nature. Condition 3 indicates the precision $\textrm{pr}(A^{*}=1\mid A=1,X=x)>0.5$, while Condition 4 states that the false omission rate $\textrm{pr}(A^{*}=1\mid A=0,X=x)<0.5$. These conditions are intuitively reasonable in many scenarios. For example, consider $A^{*}$ as the actual COVID-19 vaccination status of a patient, and $A$ as whether the patient was vaccinated at a specific hospital. If hospital records show a positive vaccination status, it must be that the patient was truly vaccinated, making $\textrm{pr}(A^{*}=1\mid A=1,X=x)=1$, thereby fulfilling Condition 2 (and Condition 1) of Assumption 4. If one can reasonably assume that the misclassification error is balanced across covariates, such that $X\!\perp\!\!\!\perp A\mid A^{*}$, i.e., the surrogate $A$ only depends on $A^{*}$, then one can safely omit $X$ from Assumption 4. We note that Hu (2008) considers alternatives to Conditions 1-4.
The informal proof given above is formalized in Theorem 1 below in a more general setting where $Y$ can be either binary, a count, or a continuous outcome, and $X$ is possibly vector-valued. For the proof of this theorem, please refer to §LABEL:supp-sec:thm_identification in the Supplementary Materials.
@@ 4 Semiparametric theory
$\Lambda^{F,\bot}=\left\{K=k\left(Y,A,A^{*},Z,X\right):E\left(K\mid Y,A^{*},X\right)=E\left(K\mid A,A^{*},X\right)=E\left(K\mid Z,A^{\ast},X\right)=0\right\}\cap L_{0}^{2,F},$ (10)
@@
The derivation of (10) and (11) is deferred to §LABEL:supp-sec:proof_tangentspace_full in the Supplementary Materials. In the following, we will make reference to the oracle nonparametric model $\mathcal{M}_{np}^{*}$ as the collection of oracle laws compatible with the DAG given in Figure 2 which places no restriction on the joint distribution of $(A^{*},Y,Z,A,X)$. We then have the following result:
@@
The efficient influence function of $\psi_{a^{*}}$ in the oracle hidden treatment model $\mathcal{M^{*}}$ is equal to the efficient influence function in the oracle nonparametric model $\mathcal{M}_{np}^{*}$.
The proof of Lemma 1 is relegated to the Supplementary Materials, §LABEL:supp-sec:eif_lemma.
@@
So far, we have characterized the tangent space, the set of influence functions and the efficient influence function for $\psi_{a^{*}}$ in $\mathcal{M^{*}}$. We now turn our attention to the observed data model $\mathcal{M}$. The following result summarizes our characterization of the tangent space of the observed data hidden treatment model $\Lambda$, and its ortho-complement, obtained by carefully applying a certain coarsening operator which maps the tangent space for the full data model, $\Lambda^{F}$, onto the $L^{2}_{0}(Y,A,Z,X)$ space.
@@
$\Lambda^{\bot}=\{L=l(Y,A,Z,X):E(L\mid Y,A^{*},X)=E(L\mid A,A^{*},X)=E(L\mid Z,A^{*},X)=0\}\cap L^{2}_{0},$
which is equivalently expressed as
@@
This implies that $h_{1}$ and $h_{2}$ belong to the null space of the $k\times 2$ matrix $M_{Y\mid A^{*},X}$. Under Assumption 2, the null space has a dimension of $k-2$ when $k\geq 2$. When $Y$ is binary, i.e., $k=2$, $M_{Y\mid A^{*},X}$ is full rank, leading to the unique solution of (13): $h_{j}(Y,X)=0,j=1,2$. This is further detailed in Corollary 1.
@@
Suppose that $A^{*},A,Z,Y$ are all binary variables, then the observed data hidden treatment model is locally nonparametric in the sense that the tangent space of $\mathcal{M}$ is $L^{2}_{0}$, i.e., $\Lambda=\{L=l(Y,A,Z,X)\in L^{2}_{0}\}$.
For more general $Y$, i.e., categorical or continuous, the following theorem provides a closed-form characterization of $\Lambda$ which facilitates influence function-based inference. In this vein, let $P_{1}=\{h(Y,X):E\{h(Y,X)\mid A^{*},X\}=0\}\cap L^{2}_{0}$ and the space of functions
$P_{2}=\Biggl{\{}v(Y,X)=\frac{v_{1}(Y,X)-E\{v_{1}(Y,X)\mid A^{*}=0,X\}}{E\{v_{1}(Y,X)\mid A^{*}=1,X\}-E\{v_{1}(Y,X)\mid A^{*}=0,X\}}$
$\qquad\qquad+\frac{v_{0}(Y,X)-E\{v_{0}(Y,X)\mid A^{*}=1,X\}}{E\{v_{0}(Y,X)\mid A^{*}=0,X\}-E\{v_{0}(Y,X)\mid A^{*}=1,X\}}-1:v_{1}(Y,X),v_{0}(Y,X)\in L^{2}\Biggr{\}}.$
Theorem 3.
Suppose that $Y$ is bounded, i.e., $|Y|<L$; then we have that $P_{1}=P_{2}$; and
$\begin{split}&\Lambda^{\bot}=\bigl{\{}\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}w_{1}(Y,X)+\{A-E(A\mid A^{*}=1,X)\}\\ &\qquad\quad\,\{Z-E(Z\mid A^{*}=0,X)\}w_{2}(Y,X):w_{1},w_{2}\in P_{2}\bigr{\}}.\end{split}$ (14)
Such equivalence provides a closed-form characterization of $P_{1}$. Based on this, we obtain a closed-form characterization of $\Lambda^{\bot}$. See §LABEL:supp-sec:orthocomplement in the Supplementary Materials for a proof of the result.
Having characterized the observed data tangent space, we note that the set of all observed data influence functions for the functional $\psi_{a^{*}}$ in the hidden treatment model can be written as $\phi+\Lambda^{\bot}$ for any influence function $\phi$. Therefore, it suffices to identify a single influence function $\phi$. To accomplish this, we establish a formal relationship between influence functions in the observed and oracle data models. As we have characterized the influence functions for the oracle model, we may leverage this relationship to derive the corresponding set of influence functions in the observed hidden treatment model. The following result provides such a characterization of observed data influence functions of $\psi_{a^{*}}$.
@@
The efficient influence function $\phi_{a^{*}}^{eff}$ is given by the orthogonal projection of $\phi_{a^{*}}$ onto the observed data tangent space of the hidden treatment model, i.e., $\phi_{a^{*}}^{eff}=\Pi(\phi_{a^{*}}\mid\Lambda)$ where $\Pi$ is the orthogonal projection operator. Therefore,
$\phi_{a^{*}}^{eff}$ $=\phi_{a^{*}}(Y,A,Z,X)-\{A-E(A\mid A^{*}=0,X)\}\{Z-E(Z\mid A^{*}=1,X)\}w^{opt}_{1}(Y,X)$
$\qquad-\{A-E(A\mid A^{*}=1,X)\}\{Z-E(Z\mid A^{*}=0,X)\}w^{opt}_{2}(Y,X).$
In the case of binary $Y$, $w^{opt}_{1}=w^{opt}_{2}=0$, and thus (4) is the efficient influence function for $\psi_{a^{*}}$ in the hidden treatment model $\mathcal{M}$. A closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ in the more general categorical outcome setting is given in the Supplemental Materials §LABEL:supp-sec:closed_form_w. As a closed-form expression for $\{w^{opt}_{j}:j=1,2\}$ and thus for the efficient influence function is not generally available for continuous $Y$, we give in the same Supplemental Materials Section a simple approach for obtaining an approximate efficient influence function based on a basis function approximation of $\{w^{opt}_{j}:j=1,2\}$.
The theorem provides an interesting insight about the relationship between an oracle influence function and the corresponding observed data influence function. Specifically, comparing (4) with the full data influence function (2), one may note that the former is obtained by substituting the indicator function $\mathbb{I}_{a^{*}}(A^{*})$ in the latter with the following expression:
@@
$=$ $\mathbb{I}_{a^{*}}(A^{*}),$
where we have used the fact that $(A,Y,Z)$ are mutually independent given $A^{*},X$ as well as the fact that
$\frac{A-E(A\mid A^{*}=1-a^{*},X)}{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)}$
and
$\frac{Z-E(Z\mid A^{*}=1-a^{*},X)}{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)}$
are conditionally unbiased for $\mathbb{I}_{a^{*}}(A^{*})$. This provides an elegant link between (2) and (4). In the full data influence function (2), the unobserved variable $A^{*}$ only appears within the indicator function $\mathbb{I}_{a^{*}}(A^{*})$. In the observed data influence function (4), $\mathbb{I}_{a^{*}}(A^{*})$ is replaced by an expression of observed variables (17), whose expectaction conditional on $(A^{*},X)$, is equal to $\mathbb{I}_{a^{*}}(A^{*})$.
@@ 5 Semiparametric estimation and inference
Typically, nuisance functions are estimated in a initial step from the observed data. However, because the nuisance functions in (18) involve the unobserved variable $A^{*}$, estimation of these functions is not straightforward, particularly using nonparametric estimators. A possible approach is to first estimate the observed data distribution, and to then compute nuisance functions following a standard plug-in principle. For a detailed discussion of this approach, please see §LABEL:supp-sec:me_estimator of the Supplementary Materials. A notable challenge of the approach is its strong dependence on the precision with which the observed data distribution $\textrm{pr}(Y,A,Z\mid X)$ can be estimated. This dependence arises because the nuisance functions are actually solutions to polynomial equations that are smooth in the observed data law $\textrm{pr}(Y,A,Z\mid X)$. This dependence renders the estimated nuisance functions excessively sensitive to the uncertainty in $\widehat{\textrm{pr}}(Y,A,Z\mid X)$, which makes it challenging to obtain estimates of the nuisance functions that are guaranteed to remain in the model. To address this, we propose an algorithm that circumvents the need to obtain an initial estimator of the observed data law. The algorithm is a semiparametric EM algorithm, inspired by the nonparametric EM algorithm proposed in Benaglia et al. (2009), see Algorithm 1 for implementation details.
@@
In practice, estimation processes are usually combined with cross-fitting to prevent over-fitting. This is achieved by using separate parts of the dataset for estimating nuisance functions and calculating the empirical mean in (18). Our algorithm is implemented using a modified version of cross-fitting. This approach slightly deviates from the standard cross-fitting procedure, as the training and evaluation sets are iterated simultaneously. Specifically, during each M-step, models $\widehat{E}(p^{(t)}\mid X),\widehat{E}(p^{(t)}\mid Y,X),\widehat{E}(p^{(t)}\mid A,X),\widehat{E}(p^{(t)}\mid Z,X)$ are developed using the training set. These models are then applied to both the training set and the evaluation set to obtain $\widetilde{\textrm{pr}}(A^{*}=1\mid X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Y_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid A_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Z_{i},X_{i})$ for $i=1,\ldots,n$. See Algorithm LABEL:supp-alg:EM_cf in the Supplementary Materials for a detailed description. It is important to note that in this process, the models evaluating the nuisances are independent of the evaluation data, making it a valid form of cross-fitting.
@@
Suppose that the nuisance functions are estimated from a separate independent sample, e.g., by cross-fitting. Moreover, suppose that Assumptions 2-5 hold, $Y$ is bounded, and
@@
Remarkably, Theorem 5 does not require nuisance functions to converge at root-$n$ rates as it suffices for their pairwise product of convergence rates to be faster than root-$n$, and for the converge rate of $\widehat{E}(A\mid a^{*}_{j},X)$ raised to the $\alpha$ power to likewise be faster than root-$n$, where $\alpha$ is a user-specified positive integer. The choice of $\alpha$ appears to offer a cost-free advantage; however, there is indeed a trade-off between $\alpha$ and the margin constant $\epsilon_{1}$ in Assumption 5. If $\epsilon_{1}$ tends to zero, $\alpha$ must also tend to zero, otherwise the constant term $\epsilon_{1}^{-\alpha}$ scaling $\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty}$ will become unbounded (see the Proof of Theorem 5 in the Supplementary Materials §LABEL:supp-sec:proof_asymptotic).